Item 6 Review Part A The electric field strength just above one face of a copper penny is 2080 N/C What is the surface charge density on this face of the penny? Express your answer with the appropriate units. alue Units Submit Request Answer Provide Feedback
Answer #1

By using Gauss law,

E*S = Q/e = (d*S) /e

d = E*e

Where, d is the charge density ,

e the electric constant and

S is an infinitesimal area parallel to the area of face of the penny

d = 2080 * 8.854*10^-12 C/m^2

= 1.84*10^-8 C/m^2